find the corresponding area between z= 0 and each of the following

1. z= 0.85

2. z= 1.27

3. z= 2.86

4. z= -1.05

5. z= -2.96


1
Expert's answer
2021-05-07T10:39:55-0400

Solution:

(a) "P(0\le z\le 0.85)=P(z\le0.85)-P(z\le0)=0.80234-0.50=0.30324"

(b)"P(0\le z\le 1.27)=P(z\le1.27)-P(z\le0)=0.89796-0.50=0.39796"

(c) "P(0\le z\le 2.86)=P(z\le2.86)-P(z\le0)=0.99788-0.50=0.49788"

(d)"P(-1.05\le z\le 0)=P(z\le0)-P(z\le-1.05)"

"=P(z\le0)-[1-P(z\le1.05)]=0.85314+0.50-1=0.35314"

(e)"P(-2.96\le z\le 0)=P(z\le0)-P(z\le-2.96)"

"=P(z\le0)-[1-P(z\le2.96)]=0.99846+0.50-1=0.49846"